# BUFFER-X: Towards Zero-Shot Point Cloud Registration in Diverse Scenes

### v1

@@ 1 Introduction
The field of deep learning-based point cloud registration has made steady and remarkable progress, including enhancing feature distinctiveness [2, 4, 5, 6, 16, 49, 32, 15], improving data association strategies [32, 91, 90, 26, 45, 46, 87, 44], and developing more robust pose estimation solvers [67, 33, 79, 94]. Consequently, existing approaches achieve strong performance on test sequences within the same dataset used for training, successfully estimating the relative pose between two partially overlapping point clouds [77, 40, 43, 78, 10, 84].
@@
More recently, there has been growing interest in tackling the generalization of these deep learning-based methods [5, 6, 54, 25, 17], which is the capability of a network to perform well across diverse real-world scenarios. While these approaches have demonstrated excellent generalization performance, in practice, most existing methods still require the user to provide optimal parameters, such as voxel size for downsampling cloud points and search radius for descriptor generation, when dealing with unseen domain datasets. In this paper, we refer to this manual tuning as an oracle. Therefore, it is still desirable to develop zero-shot registration approaches for better usability and practical deployment.
@@
Fig. 1: Sucess rate (unit: %) of zero-shot point cloud registration with state-of-the-art approaches on 11 datasets [28, 88, 34, 32, 81, 60, 61, 69, 71, 56]. Without any prior information or manual parameter tuning for the test datasets, our BUFFER-X shows robust generalization capability across diverse scenes even though the network is only trained on the 3DMatch dataset [88].
@@
Furthermore, despite nearly a decade of research in deep learning-based registration, most studies remain confined to specific scenarios, primarily conducting experiments using omnidirectional LiDAR point clouds for outdoor environments [82, 28] and RGB-D depth clouds for indoor settings [88]. For this reason, domain generalization experiments on LiDAR point clouds in indoors [60] or with different LiDAR scan patterns in outdoor environments [60, 34, 39] are less explored. This underscores the need for a new benchmark that better reflects real-world sensor variations to evaluate generalizability across unseen environments and diverse scanning patterns.
@@
In this context, the main contribution of this paper is addressing two key issues above and propose: a) a zero-shot registration architecture and b) a novel benchmark to help evaluate the generalization capability of deep learning-based registration approaches, as shown in Fig. 1. First, inspired by the remarkable generalization of BUFFER [5] as long as a user manually tunes the voxel size and search radius, we first thoroughly analyze the architectural principles that underpin its generalization. Then, we identify three factors that hinder the zero-shot capabilities of existing methods in Sec. 3. Building on these insights, we introduce a self-adaptive mechanism to determine the optimal voxel size for each test scene and streamline the pipeline of BUFFER, ultimately presenting a robust multi-scale patch-wise approach. We name our approach BUFFER-X to signify that it is an extension of BUFFER.
@@ 2 Related Work
3D point cloud registration, which estimates the relative pose between two partially overlapping point clouds, is a fundamental problem in the fields of robotics and computer vision [7, 83, 18, 13, 37]. Overall, point cloud registration methods are classified into two categories based on whether their performance relies on the availability of an initial guess for registration: a) local registration [11, 66, 55, 36, 57, 72] and b) global registration [27, 24, 78, 94, 10, 76, 77, 43, 40]. Global registration methods can be further classified into two types: a) correspondence-free [62, 12, 10, 52, 23, 19, 65, 50, 31, 51, 93] and b) correspondence-based [27, 78, 94, 24, 38, 87, 44, 86] approaches. In this study, we focus on the latter and particularly place more emphasis on deep learning-based registration methods.
@@
Since Qi et al. [58] demonstrated that learning-based techniques in 2D images can also be applied to 3D point clouds, a wide range of learning-based point cloud registration approaches have been proposed. Building on these advances, novel network architectures with increased capacity have continuously emerged, ranging from MinkUNet [22, 20, 21], cylindrical convolutional network [6, 5, 95], KPConv [32, 70, 8] to Transformers [59, 74, 73, 16].
@@
While these advancements have led to improved registration performance, some of these methods often exhibit limited generalization capability, leading to performance degradation when applied to point clouds captured by different sensor configurations or in unseen environments. To tackle the generalization problem, Ao et al. [6, 3] introduce SpinNet, a patch-based method that normalizes the range of local point coordinates within a fixed-radius neighborhood to $[-1,1]$. This makes us come to realize that patch-wise scale normalization is key to achieving a data-agnostic registration pipeline.
@@
Further, Ao et al. [5] proposed BUFFER to enhance efficiency by combining point-wise feature extraction with patch-wise descriptor generation. However, we found that such learning-based keypoint detectors can hinder robust generalization, as their failure in out-of-domain distributions may trigger a cascading failure in subsequent steps; see Sec. 3.2.3. In addition, despite the high generalizability of BUFFER, we observed that during cross-domain testing, users had to manually specify the optimal voxel size for the test data, which hinders fully zero-shot inference.
@@ 3.1 Problem statement
Fig. 2: (a) Variation in the number of points after voxelization at different voxel sizes $v$ for multiple datasets. Note that even in indoor environments, the number of points varies significantly depending on the sensor (i.e.,TIERS [60] from three different LiDAR sensors vs. 3DMatch [88] from an RGB-D camera). Additionally, even both TIERS and KITTI are acquired using the same sensor type (i.e., omnidirectional LiDAR sensors), the number of voxel-sampled points can differ greatly depending on the environment (i.e.,TIERS, collected in indoor scenes vs. KITTI [28], acquired in outdoor environments using a 64-channel LiDAR sensor). (b) Empirical distribution of the datasets’ maximum range.
@@ 3.2.1 Voxel size and search radius
First, dependencies of optimal search radius $r$ for local descriptors and voxel size $v$ for each dataset are problematic. The optimal parameters vary significantly across datasets due to differences in scale and point density (e.g., small indoor scenes vs. large outdoor spaces [5]) Consequently, improper $r$ or $v$ can severely degrade registration performance by failing to account for specific scale and density characteristics of a given environment or sensor; see Sec. 5.1. For instance, in Fig. 2(a), as $v$ controls the maximum number of points that can be fed into the network, a too-small $v$ can trigger out-of-memory errors when outdoor data processed with parameters optimized for indoor environments are taken as input to the network.
@@ 3.2.2 Input scale normalization
Next, directly feeding raw $x$, $y$, and $z$ values into the network leads to strong in-domain dependency [32, 59]. That is, when a model fits to the training distribution, large scale discrepancies between training and unseen data can cause catastrophic failure (see Fig. 2(b) for an example of maximum range discrepancy). For this reason, we conclude that normalizing input points within local neighborhoods (or patches) is necessary to achieve generalizability, ensuring that their coordinates lie within a bounded range (e.g., $[-1,1]$[6, 5].
@@ 3.2.3 Keypoint detection
Following Sec. 3.2.2, we observed that point-wise feature extractor modules in existing methods [32, 59, 80] are empirically brittle to out-of-domain data. Because failed keypoint detection leads to the selection of unreliable and non-repeatable points as keypoints, it results in low-quality descriptors and ultimately degrades the quality of $\mathcal{A}$ [30].
@@ 4 BUFFER-X
Building upon our key observations in Sec. 3.2, we present our multi-scale zero-shot registration pipeline; see Fig. 3. First, the appropriate voxel size and radii for each cloud pair are predicted by geometric bootstrapping (Sec. 4.1), considering the overall distribution of cloud points and the density of neighboring points, respectively. Then, we extract Mini-SpinNet-based features [5] for the sampled points via FPS at multiple scales (Sec. 4.2). Finally, at the intra- and cross-scale levels, refined correspondences are estimated based on consensus maximization [68, 67, 89] (Sec. 4.3) and serve as input for the final relative pose estimation using a solver.
Fig. 3: Overview of our BUFFER-X, which mainly consists of three steps. (a) Geometric bootstrapping (Sec. 4.1) to determine the appropriate voxel size and radii for the given source $\mathcal{P}$ and target $\mathcal{Q}$ clouds. (b) Multi-scale patch embedder (Sec. 4.2) to generate patch-wise descriptor $\mathcal{S}_{\xi}$ for multiple scale $\xi\in\{l,m,g\}$, where $l$, $m$, and $g$ represent local, middle, and global scales, respectively. Specifically, Mini-SpinNet [5] outputs cylindrical feature maps $\mathcal{C}_{\xi}$ and vector feature set $\mathcal{F}_{\xi}$. (c) Hierarchical inlier search (Sec. 4.3), which first performs nearest neighbor-based intra-scale matching using $\mathcal{F}^{\mathcal{P}}_{\xi}$ and $\mathcal{F}^{\mathcal{Q}}_{\xi}$ at each scale, followed by pairwise transformation estimation. Finally, it identifies globally consistent inliers $\mathcal{I}$ across all scales to refine correspondences based on consensus maximization [68, 89].
@@ 4.1 Geometric bootstrapping
Sphericity-based voxelization. First, we determine proper voxel size $v$ by leveraging sphericity, quantified using eigenvalues [29, 1], to reflect how the cloud points are dispersed in space. To this end, we apply principal component analysis (PCA) [42] to the covariance of sampled points, which can efficiently capture point dispersion by analyzing eigenvalues while remaining computationally lightweight.
@@
which are assumed to be $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$. Then, using these properties, we can compute the sphericity $\frac{\lambda_{3}}{\lambda_{1}}$ [1], which quantifies how evenly a point cloud is distributed in space. Since LiDAR points are primarily distributed along the sensor’s horizontal plane (i.e., forming a disc-like shape), $\frac{\lambda_{3}}{\lambda_{1}}$ tends to be low compared to RGB-D point clouds.
@@
Hereafter, for brevity, we denote $f_{v}(\mathcal{P})$ and $f_{v}(\mathcal{Q})$ simply as $\mathcal{P}$ and $\mathcal{Q}$, respectively,
Density-aware radius estimation. Next, in contrast to some state-of-the-art approaches [5, 6] that use a single fixed user-defined search radius, we determine $r$ at local, middle, and global scales, respectively, by considering the input point densities. Let neighboring search function within the radius $r$ around a query point ${\bm{p}}_{q}$ be:
@@ 4.2 Multi-scale patch embedder
Farthest point sampling. As discussed in Sec. 3.2.3, we sample ${\mathcal{P}_{\xi}}$ from $\mathcal{P}$ at each scale using FPS to be free from a learning-based keypoint detector (resp. ${\mathcal{Q}_{\xi}}$ from $\mathcal{Q}$). Note that instead of extracting local, middle, and global-scale descriptors for the same sampled point [85], we independently sample separate points for each scale, as illustrated in Fig. 3(b). This is because we empirically found that different regions may require distinct scales for optimal feature extraction; see Sec. 5.3.
Mini-SpinNet-based descriptor generation. Using multiple radii $r_{\xi}$, we sample patches at three distinct scales, providing a more comprehensive multi-scale representation. Then, we use Mini-SpinNet[5] for descriptor generation, which is a lightweight version of SpinNet [6].
In particular, building on the insights from Sec. 3.2.2, we ensure the scale of points in each patch is normalized to a bounded range of $[-1,1]$ by dividing by $r_{\xi}$; see Fig. 4(b). By doing so, we can resolve the dependency on the in-domain scale. To maintain consistency across patches at all scales, we fix the patch size to $N_{\text{patch}}$ and randomly sample when a patch exceeds this size, ensuring a consistent number of points regardless of scale variations.
Fig. 4: (a) Visual description of local ($r_{l}$), middle ($r_{m}$), and global ($r_{g}$) radii for the same point to illustrate scale differences and (b) normalized patches ranging from $[-1,1]$. Note that their reference frames follow the eigenvectors obtained from principal component analysis (PCA) [42, 1]. The $z$-axis is assigned to the eigenvector ${\bm{v}}_{3}$, which corresponds to the smallest eigenvalue.
Finally, taking these normalized patches as inputs, Mini-SpinNet outputs a superset $\mathcal{S}^{\mathcal{P}}_{\xi}$ consisting of $D$-dimensional feature vectors $\mathcal{F}^{\mathcal{P}}_{\xi}$ and cylindrical feature maps $\mathcal{C}^{\mathcal{P}}_{\xi}$, where corresponds to ${\mathcal{P}_{\xi}}$ (resp. $\mathcal{S}^{\mathcal{Q}}_{\xi}$ consisting of $\mathcal{F}^{\mathcal{Q}}_{\xi}$ and $\mathcal{C}^{\mathcal{Q}}_{\xi}$ from ${\mathcal{Q}_{\xi}}$), as described in Fig. 3. Note that while BUFFER [5] utilizes learned reference axes to extract cylindrical coordinates, our approach defines the reference axes for each patch by applying PCA to the covariance of points within the patch, setting the $z$-direction as ${\bm{v}}_{3}$ (as in 2 and illustrated by $z={\bm{v}}_{3}$ in Fig. 4(b)), to eliminate potential dataset-specific inductive biases.

### v3

@@ 1 Introduction
The field of deep learning-based point cloud registration has made steady and remarkable progress, including enhancing feature distinctiveness [Ao20ietcv-SGHs, Ao20pr-Repeatable, Ao23CVPR-BUFFER, Ao21cvpr-Spinnet, Chen24cpvr-DCATr, Mu24cvpr-ColorPCR, Huang21cvpr-PREDATORRegistration, Chen24icra-Tree-based-Transformer], improving data association strategies [Huang21cvpr-PREDATORRegistration, Zhang24cvpr-FastMAC, Zhang23cvpr-MaxCliques, Fathian24ral-Clipperplus, Liu24tgrs-DeepSemanticMatching, Liu24tpami-NCMNet, Yuan24cvpr-InlierConfidence, Liu24cvpr-Extend], and developing more robust pose estimation solvers [Shi24ral-RANSAC, Huang24cvpr-Scalable, Yang24tpami-MAC, Zhou16eccv-FastGlobalRegistration]. Consequently, existing approaches achieve strong performance on test sequences within the same dataset used for training, successfully estimating the relative pose between two partially overlapping point clouds [Yang20tro-teaser, Lim24ijrr-Quatropp, Lim22icra-Quatro, Yang16pami-goicp, Bernreiter21ral-PHASER, Yin23icra-Segregator].
More recently, there has been growing interest in tackling the generalization of these deep learning-based methods [Ao23CVPR-BUFFER, Ao21cvpr-Spinnet, Poiesi22pami-GeDi, Dosovitskiy19iclr-YOTO, Chen21nips-OnlyTrainOnce], which is the capability of a network to perform well across diverse real-world scenarios.
@@
While these approaches have demonstrated excellent generalization performance, in practice, most existing methods still require the user to provide optimal parameters, such as voxel size for downsampling cloud points and search radius for descriptor generation, when dealing with unseen domain datasets. In this paper, we refer to this manual tuning as an oracle. Therefore, it is still desirable to develop zero-shot registration approaches for better usability and practical deployment.
@@
Fig. 1: Success rate (unit: %) of zero-shot point cloud registration with state-of-the-art approaches on 11 datasets [Geiger13ijrr-KITTI, Zeng17cvpr-3dmatch, Jung23ijrr-HeLiPR, Huang21cvpr-PREDATORRegistration, Yeshwanth23iccv-Scannet++, Qingqing22iros-TIERS, Ramezani20iros-NewerCollege, Sun20cvpr-WaymoDataset, Tian23iros-KimeraMultiExperiments, Pomerleau12ijrr-ETH]. Without any prior information or manual parameter tuning for the test datasets, our BUFFER-X shows robust generalization capability across diverse scenes even though the network is only trained on the 3DMatch dataset [Zeng17cvpr-3dmatch].
@@
Furthermore, despite nearly a decade of research in deep learning-based registration, most studies remain confined to specific scenarios, primarily conducting experiments using omnidirectional LiDAR point clouds for outdoor environments [Yew18eccv-3dfeatnet, Geiger13ijrr-KITTI] and RGB-D depth clouds for indoor settings [Zeng17cvpr-3dmatch]. For this reason, domain generalization experiments on LiDAR point clouds in indoors [Qingqing22iros-TIERS] or with different LiDAR scan patterns in outdoor environments [Qingqing22iros-TIERS, Jung23ijrr-HeLiPR, Lim24iros-HeLiMOS] are less explored. This underscores the need for a new benchmark that better reflects real-world sensor variations to evaluate generalizability across unseen environments and diverse scanning patterns.
@@
In this context, the main contribution of this paper is addressing two key issues above and propose: a) a zero-shot registration architecture and b) a novel benchmark to help evaluate the generalization capability of deep learning-based registration approaches, as shown in Fig. 1. First, inspired by the remarkable generalization of BUFFER [Ao23CVPR-BUFFER] as long as a user manually tunes the voxel size and search radius, we first thoroughly analyze the architectural principles that underpin its generalization. Then, we identify three factors that hinder the zero-shot capabilities of existing methods in Sec. 3. Building on these insights, we introduce a self-adaptive mechanism to determine the optimal voxel size for each test scene and streamline the pipeline of BUFFER, ultimately presenting a robust multi-scale patch-wise approach. We name our approach BUFFER-X to signify that it is an extension of BUFFER.
@@ 2 Related Work
3D point cloud registration, which estimates the relative pose between two partially overlapping point clouds, is a fundamental problem in the fields of robotics and computer vision [Aoki24icra-3DBBS, Yin24ijcv-LiDARLocSurvey, Chen22ar-Overlapnet, Cattaneo22tro-LCDNet, Lee22arxiv-LearningReg]. Overall, point cloud registration methods are classified into two categories based on whether their performance relies on the availability of an initial guess for registration: a) local registration [Besl92pami, Segal09rss-GeneralizedICP, Pomerleau13auro-ICPcomparison, Koide21icra-VGICP, Oomerleau12ijrr-ethpc, Vizzo23ral-KISSICP] and b) global registration [Fischler81, Dong17isprsremotesensing-GHICP, Yang16pami-goicp, Zhou16eccv-FastGlobalRegistration, Bernreiter21ral-PHASER, Yang19rss-teaser, Yang20tro-teaser, Lim22icra-Quatro, Lim24ijrr-Quatropp]. Global registration methods can be further classified into two types: a) correspondence-free [Rouhani11iccv-CorrespondenceFreeReg, Brown19pr-AFamiliyofBnB, Bernreiter21ral-PHASER, Papazov12ijrr-Rigid3DGeometryMatching, Chum03jprs-LocallyOptimizedRANSAC, Choi97jcv-RANSAC, Schnabel07cgf-EfficientRANSAC, Olsson09pami-bnbRegistration, Hartley09ijcv-globalRotationRegistration, Pan19robotbiomim-MultiViewBnB, Zhao24cvpr-CorrespondenceFree] and b) correspondence-based [Fischler81, Yang16pami-goicp, Zhou16eccv-FastGlobalRegistration, Dong17isprsremotesensing-GHICP, Lei17tip-FastDescriptors, Yuan24cvpr-InlierConfidence, Liu24cvpr-Extend, Yu24cvpr-InstanceAware] approaches. In this study, we focus on the latter and particularly place more emphasis on deep learning-based registration methods.
@@
Since Qi et al. [Qi17cvpr-pointnet] demonstrated that learning-based techniques in 2D images can also be applied to 3D point clouds, a wide range of learning-based point cloud registration approaches have been proposed. Building on these advances, novel network architectures with increased capacity have continuously emerged, ranging from MinkUNet [Choi19iccv-FCGF, Choy20cvpr-deepGlobalRegistration, choy19cvpr-4DSTConv], cylindrical convolutional network [Ao21cvpr-Spinnet, Ao23CVPR-BUFFER, Zhu21cvpr-Cylindrical3D], KPConv [Huang21cvpr-PREDATORRegistration, Thomas19iccv-kpconv, Bai20cvpr-D3Feat] to Transformers [Qin23tpami-GeoTransformer, Wu22nips-PointTransformerV2, Wu24cvpr-PointTransformerV3, Chen24cpvr-DCATr].
@@
While these advancements have led to improved registration performance, some of these methods often exhibit limited generalization capability, leading to performance degradation when applied to point clouds captured by different sensor configurations or in unseen environments. To tackle the generalization problem, Ao et al. [Ao21cvpr-Spinnet, ao22tpami-YOTO] introduced SpinNet, a patch-based method that normalizes the range of local point coordinates within a fixed-radius neighborhood to [1,1][-1,1]. This makes us come to realize that patch-wise scale normalization is key to achieving a data-agnostic registration pipeline.
@@
Further, Ao et al. [Ao23CVPR-BUFFER] proposed BUFFER to enhance efficiency by combining point-wise feature extraction with patch-wise descriptor generation. However, we found that such learning-based keypoint detectors can hinder robust generalization, as their failure in out-of-domain distributions may trigger a cascading failure in subsequent steps; see Sec. 3.2.3. In addition, despite the high generalizability of BUFFER, we observed that during cross-domain testing, users had to manually specify the optimal voxel size for the test data, which hinders fully zero-shot inference.
@@ 3.1 Problem statement
Fig. 2: (a) Variation in the number of points after voxelization with different voxel sizes $v$ across datasets. Even in indoor scenes, point counts vary significantly depending on the sensor type (i.e.,TIERS [Qingqing22iros-TIERS] vs. 3DMatch [Zeng17cvpr-3dmatch]). Notably, TIERS and KITTI [Geiger13ijrr-KITTI], both using omnidirectional LiDARs, yield different point densities due to indoor vs. outdoor environments. (b) Empirical distribution of the datasets’ maximum range.
@@ 3.2.1 Voxel size and search radius
First, dependencies of optimal search radius $r$ for local descriptors and voxel size $v$ for each dataset are problematic. The optimal parameters vary significantly across datasets due to differences in scale and point density (e.g., small indoor scenes vs. large outdoor spaces [Ao23CVPR-BUFFER]) Consequently, improper $r$ or $v$ can severely degrade registration performance by failing to account for specific scale and density characteristics of a given environment or sensor; see Sec. 5.1. For instance, in Fig. 2(a), as $v$ controls the maximum number of points that can be fed into the network, a too-small $v$ can trigger out-of-memory errors when outdoor data processed with parameters optimized for indoor environments are taken as input to the network.
@@ 3.2.2 Input scale normalization
Next, directly feeding raw $x$, $y$, and $z$ values into the network leads to strong in-domain dependency [Huang21cvpr-PREDATORRegistration, Qin23tpami-GeoTransformer]. That is, when a model fits to the training distribution, large scale discrepancies between training and unseen data can cause catastrophic failure (see Fig. 2(b) for an example of maximum range discrepancy). For this reason, we conclude that normalizing input points within local neighborhoods (or patches) is necessary to achieve generalizability, ensuring that their coordinates lie within a bounded range (e.g., [1,1][-1,1][Ao21cvpr-Spinnet, Ao23CVPR-BUFFER].
@@ 3.2.3 Keypoint detection
Following Sec. 3.2.2, we observed that point-wise feature extractor modules in existing methods [Huang21cvpr-PREDATORRegistration, Qin23tpami-GeoTransformer, Yao24iccv-PARENet] are empirically brittle to out-of-domain data. Because failed keypoint detection leads to the selection of unreliable and non-repeatable points as keypoints, it results in low-quality descriptors and ultimately degrades the quality of $\mathcal{A}$ [Harris88avc-HarrisCorner].
@@ 4 BUFFER-X
Building upon our key observations in Sec. 3.2, we present our multi-scale zero-shot registration pipeline; see Fig. 3. First, the appropriate voxel size and radii for each cloud pair are predicted by geometric bootstrapping (Sec. 4.1), considering the overall distribution of cloud points and the density of neighboring points, respectively. Then, we extract Mini-SpinNet-based features [Ao23CVPR-BUFFER] for the sampled points via FPS at multiple scales (Sec. 4.2). Finally, at the intra- and cross-scale levels, refined correspondences are estimated based on consensus maximization [Sun22ral-TriVoC, Shi24ral-RANSAC, Zhang24tpami-AcceleratingGloballyCM] (Sec. 4.3) and serve as input for the final relative pose estimation using a solver.
Fig. 3: Overview of our BUFFER-X, which mainly consists of three steps. (a) Geometric bootstrapping (Sec. 4.1) to determine the appropriate voxel size and radii for the given source $\mathcal{P}$ and target $\mathcal{Q}$ clouds. (b) Multi-scale patch embedder (Sec. 4.2) to generate patch-wise descriptor $\mathcal{S}_{\xi}$ for multiple scale $\xi\in\{l,m,g\}$, where $l$, $m$, and $g$ represent local, middle, and global scales, respectively. Specifically, Mini-SpinNet [Ao23CVPR-BUFFER] outputs cylindrical feature maps $\mathcal{C}_{\xi}$ and vector feature set $\mathcal{F}_{\xi}$. (c) Hierarchical inlier search (Sec. 4.3), which first performs nearest neighbor-based intra-scale matching using $\mathcal{F}^{\mathcal{P}}_{\xi}$ and $\mathcal{F}^{\mathcal{Q}}_{\xi}$ at each scale, followed by pairwise transformation estimation. Finally, it identifies globally consistent inliers $\mathcal{I}$ across all scales to refine correspondences based on consensus maximization [Sun22ral-TriVoC, Zhang24tpami-AcceleratingGloballyCM].
@@ 4.1 Geometric bootstrapping
Sphericity-based voxelization. First, we determine the proper voxel size $v$ by leveraging sphericity, quantified using eigenvalues [Hansen21remotesensing-Classification, Alexiou24jivp-PointPCA], to reflect how the cloud points are dispersed in space. To this end, we apply principal component analysis (PCA) [Lim21ral-Patchwork] to the covariance of sampled points, which can efficiently capture point dispersion by analyzing eigenvalues while remaining computationally lightweight.
@@
which are assumed to be $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$. Then, using these properties, we can compute the sphericity $\frac{\lambda_{3}}{\lambda_{1}}$ [Alexiou24jivp-PointPCA], which quantifies how evenly a point cloud is distributed in space. Since LiDAR points are primarily distributed along the sensor’s horizontal plane (i.e., forming a disc-like shape), $\frac{\lambda_{3}}{\lambda_{1}}$ tends to be low compared to RGB-D point clouds.
@@
Hereafter, for brevity, we denote $f_{v}(\mathcal{P})$ and $f_{v}(\mathcal{Q})$ simply as $\mathcal{P}$ and $\mathcal{Q}$, respectively.
Density-aware radius estimation. Next, in contrast to some state-of-the-art approaches [Ao23CVPR-BUFFER, Ao21cvpr-Spinnet] that use a single fixed user-defined search radius, we determine $r$ at local, middle, and global scales, respectively, by considering the input point densities. Let neighboring search function within the radius $r$ around a query point ${\bm{p}}_{q}$ be:
@@ 4.2 Multi-scale patch embedder
Farthest point sampling. As discussed in Sec. 3.2.3, we sample $\mathcal{P}_{\xi}$} from $\mathcal{P}$ at each scale using FPS to be free from a learning-based keypoint detector (resp. $\mathcal{Q}_{\xi}$} from $\mathcal{Q}$). Note that instead of extracting local, middle, and global-scale descriptors for the same sampled point [Yu21nips-CofiNet], we independently sample separate points for each scale, as illustrated in Fig. 3(b). This is because we empirically found that different regions may require distinct scales for optimal feature extraction; see Sec. 5.3.
Mini-SpinNet-based descriptor generation. Using multiple radii $r_{\xi}$, we sample patches at three distinct scales, providing a more comprehensive multi-scale representation. Then, we use Mini-SpinNet[Ao23CVPR-BUFFER] for descriptor generation, which is a lightweight version of SpinNet [Ao21cvpr-Spinnet].
In particular, building on the insights from Sec. 3.2.2, we ensure the scale of points in each patch is normalized to a bounded range of [1,1][-1,1] by dividing by $r_{\xi}$; see Fig. 4(b). By doing so, we can resolve the dependency on the in-domain scale. To maintain consistency across patches at all scales, we fix the patch size to $N_{\text{patch}}$ and randomly sample when a patch exceeds this size, ensuring a consistent number of points regardless of scale variations.
Fig. 4: (a) Visual description of local ($r_{l}$), middle ($r_{m}$), and global ($r_{g}$) radii for the same point to illustrate scale differences and (b) normalized patches ranging from [1,1][-1,1]. Note that their reference frames follow the eigenvectors obtained from principal component analysis (PCA) [Lim21ral-Patchwork, Alexiou24jivp-PointPCA]. The $z$-axis is assigned to the eigenvector ${\bm{v}}_{3}$, which corresponds to the smallest eigenvalue.
Finally, taking these normalized patches as inputs, Mini-SpinNet outputs a superset $\mathcal{S}^{\mathcal{P}}_{\xi}$ consisting of $D$-dimensional feature vectors $\mathcal{F}^{\mathcal{P}}_{\xi}$ and cylindrical feature maps $\mathcal{C}^{\mathcal{P}}_{\xi}$, where corresponds to $\mathcal{P}_{\xi}$} (resp. $\mathcal{S}^{\mathcal{Q}}_{\xi}$ consisting of $\mathcal{F}^{\mathcal{Q}}_{\xi}$ and $\mathcal{C}^{\mathcal{Q}}_{\xi}$ from $\mathcal{Q}_{\xi}$}), as described in Fig. 3. Note that while BUFFER [Ao23CVPR-BUFFER] utilizes learned reference axes to extract cylindrical coordinates, our approach defines the reference axes for each patch by applying PCA to the covariance of points within the patch, setting the $z$-direction as ${\bm{v}}_{3}$ (as in 2 and illustrated by $z={\bm{v}}_{3}$ in Fig. 4(b)), to eliminate potential dataset-specific inductive biases.